Apr 16, 2020 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

  1. Check switches on FCAL board with high scaler rates and make sure it is set to 2 V full scale. (Colin and Sasha)

Long term items

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run Updates
    1. FCAL
    2. BCAL
  4. Calibrations
    • Batch 1: 71350 - 71591 (150, 250 nA)
    • Batch 2: 71592 - 71727 (350 nA)
    1. Gain calibration (Karthik) - GainsSpring2020B2
    2. Nonlinearity corrections (Karthik) - Plots shown are after applying nonlinear corrections NonlinearSpring2020B2
    • Batch 3: 71728 - 71835 (350 nA)
    • PrimEx Calibration PrimExupdate
  5. Monitoring
    1. BCAL-LED update (Varun)[2]
    2. Summary of batch 1 2019-11 run - BCAL (Tolga) [3]
  6. Reconstruction
  7. Efficiencies
  8. Simulations
  9. NIM articles
  10. Any other business


Attending: Colin, Mark, Tolga, Varun, Karthik, Zisis, Curtis, Sean, Elton, Matt, Igal, Christina, Jon, Susan, Sasha, Chandra"

  1. Announcements
    1. Elton announced that he will be entering phased retirement, starting in August. Zisis will take over the CALWG Meetings starting today and it is likely that Mark will take over Elton's duties in Hall D, in due time.
    2. Sean mentioned that the monitoring launch for spring 2020 Batch 1 went well. Tolga looked at the BCAL and Susan ran checks. On Friday, Batch 2 and 3 will be launched, provided nobody puts a stop to that.
  2. Action Items
    1. Colin said that the action item above will be dealt with once MEDCON restrictions have been lifted.
  3. Run Updates
    1. FCAL: none
    2. BCAL: none
  4. Calibrations
    1. Gain calibration (Karthik): After 3 iterations the pi0 calibration of Batch 2 (B2) converged to a mass of 0.1346 GeV/c2 and sigma/mean = 0.0580. The gains are a bit different from B1; likely not a significant difference. The B2 gains will be applied to B3 and we will see if there is any rate dependence (B1- 150 vs B2,B3-250 nA). Mark noted tahat during spring 2020 the FADC temps were not fixered so those drifted and hot channels di took following atmospheric conditions over the duration of the spring run. Because of the survey ball incident on the BCAL at Module 26, it was decided to spring Batch 4 into runs before the incident (up to 71860) and after. Sean made this change right after the meeting. Mark had seen the U/D ratios change after that incident. Karthik will proceed to B3 on Monday (working on PrimEx until then).
      • Batch 1: 71350 - 71591 (150, 250 nA); Batch 2: 71592 - 71727 (350 nA); Batch 3: 71728 - 71835 (350 nA)
    2. Nonlinearity corrections (Karthik): The non-linear correction seems to work well (see slide 3 from above link). Karthik will follow up by repeating the study of fixing the anchor point (currently with a value of 0.95 at 11 GeV).
    3. PrimEx Calibration (Karthik): Karthik circulated an email yesterday, which is linked to above. It showed the Layer3 sums, and outliers to the fits. It converged to a mass of 0.1346, with Layers 1 and too being tight around that value and L3 more spread out.
    4. FCAL Calibration (Colin): The B1 calibrations should in principle work for the subsequent batches but this will be checked.
  5. Monitoring
    1. BCAL-LED update (Varun): LED plots for the entire spring 2020 period were shown. These show a slope over time, which is more pronounced for the downstream SiPMs. The slopes were extracted and plotted against channel number: DS SiPMs are more clustered, US very spread out. Many fits failed in a particular way; these will be investigated. Matt asked the question: does channel 400 loop back to zero, showing continuity? Mark: some of these features were not seen in the U/D ratios and may be connected to what Karthik is seeing. Varun will break things down in batches, using the run average over the entire period as his reference. If a comparison to Karthik's results is sought, then layers have to be averaged in the LED data. The slopes on slide 6 will be fit and the failed fits rechecked. Module 26 was shown and the hit is clearly visible. Mark had also observed the effect in adjacent modules, up to 3 on either side of Module 26. Channel 118 was shown; this might have been dealt with before. It needs to be rechecked. Zisis proposed tracking all abnormal channels from here on, including past history. Sean and Zisis will coordinate to set this up on a Wiki page, with links to blog entries. Example of LED fluctuations were shown; some of these are very small and typically we do not worry about them, especially L4.
    2. Summary of batch 1 Monitoring 2019-11 run - BCAL (Tolga): Tolga summarized his close examination of B1 monitoring histos for the BCAL (Susan also checked them). His detailed comments were entered in Sean's Google Docs table. Occupancy and efficiency look mostly ok. On slide 5, Δt vs BCAL Neutral Shower vertical bands appear, parallel to the y-axis; approximately 30 out of 115 plots showed this behaviour. The master trigger factory should be checked to see why these are leaking in. On slide 7, bottom right plot, a dip is seen at 70 cm. This should be examined further.
  6. Reconstruction: nothing to report.
  7. Efficiencies: nothing to report.
  8. Simulations
    • Pi0 mass in BCAL: Jon showed a pi0 gun study using Geant4 (polar angle 15-35 deg, large momentum range). Pi0 mass is about 2.5 MeV lower when magnetic field is off. This was done using MCWrapper and two specific run numbers (~30000 GlueX, ~60000 PrimEx). A discussion ensued, and several checks should be made: is tracking used and is vertex fixed (Igal), check BCAL Inv Mass module (Jon), run monoenergetic photon gun with field ON/OFF and inspect BCAL objects sigrow, sigtheta, sigphi to see if there is broadening (Matt). Matt noted that the original clusterizer clusters in solid angle, so there is a weak dependence on the target vertex. Karthik noted that GlueX fall 2018 with fixed vertex gave pi0 at 0.1335 initially and PrimEx at 0.130. Jon will check these.
  9. NIM articles: nothing to report; item will be removed from standard agenda.
  10. Any other business: Karthik asked if he should upload constants to CCDB. Elton and Sean said yes.